OpenStudy (anonymous):

What is the degree of each monomial? 7m^6n^5

OpenStudy (anonymous):

A. 5 B. 11 C. 6 D. 7

OpenStudy (anonymous):

hi

OpenStudy (anonymous):

i say b oh hi there

OpenStudy (anonymous):

what is "degree" mean?? what do you think!

OpenStudy (anonymous):

honestly dont know

OpenStudy (anonymous):

if you know that than you know the answer!

OpenStudy (anonymous):

huh??

OpenStudy (anonymous):

can u show me how to do this ??

OpenStudy (anonymous):

"A monomial is a number, a variable or a product of a number and a variable where all exponents are whole numbers".

OpenStudy (anonymous):

all you have to do is google it!!

OpenStudy (anonymous):

its c

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

The degree of the monomial is the sum of the exponents of all included variables. Constants have the monomial degree of 0. If we look at our examples above we can see that 42, 5x , 14x^12 , 2pq Monomial Degree 42 0 5x 0 + 1 = 1 14x12 0 + 12 = 12 2pq 0 + 1 + 1 = 2

OpenStudy (anonymous):

I still don't get it...